These are the news starting 01 Jan 2011
Click here for the news of the previous years.
|23 Dec 11||Recalibration of the repaired AC31M.: there remains a serious problem with the NO2 calibration factor, so NO2 readings should be taken with circonspection.|
|The AC31M NO/NO2 sensor has beeen repaired by TSE. A further calibration will be done tomorrow.|
|21 Dec 11||Maintenance and recalibration work of the CO2 and O3 sensors.|
|11 Dec 11||November 2011 data uploaded to data archive.|
|14 Nov 11||
The ultrasonic anemometer became bad yesterday 13th Nov. during the
afternoon. Please use channel 61 = cup anemometer readings as a replacement.
The windchill/wind-speed graphs will show these cup anemometers readings
from now on, until the ultrasonic anemometer will be repaired.|
Closer inspection shows that the ultrasonic anemometer became (again!) a victim of the *¨**x%& crows: a bloody bird tore one of the 4 US-sensors out of its fixation, so that it is now dangling down. I will not have the time to dismount the US-anemoter during the next weeks (this involves some scary climbing); so the cup anemomter will have to suffice. I made a quick compare over the last 8 days between both: their is a small offset between the higher mounted US-anemometer and the cup anemometer, which is 0.38 m/s lower . The linearity is excellent: see graph here!
|05 Nov 11||Maintenance and calibration work.|
|02 Nov 11||October 2011 data uploaded to data archive.|
|21 Oct 11||"Repaired" pump didn't survive very long. I replaced the old pump with a spare one (an used model, refurbished with a new ball bearing). Hopefully this one will last some time....|
|18 Oct 11||This seems to be the weekend w.e.g.w. (where everything goes wrong): the pump of the ozone sensor blocked during the 14th October (a ball bearing is terminally ill). I made a quick repair this morning, but the pump must be replaced asap.|
|15 Oct 11||Ultrasonic windsensor problem solved. Cause too silly to mention without loosing face...|
|14 Oct 11||
1. Maintenance by Envicontrol. The NOx sensor has been repaired (a nasty
problem with the molybdenum oven which converts NO into NO2, and which was
blocked by corrosion).|
2. The ultrasonic anemometer is out of order, but the mechanical one (ch.61) is working fine. If I am unable to make a quick repair tomorrow, the mechanical data will be used asap for the live data display and the windchill.
|01 Oct 11||September 2011 data uploaded to data archive.|
|21 Sep 11||August 2011 data uploaded to data archive.|
|12 Sep 11||I have added a new link to a page with energy widgets showing the yearly world cumulative energy consumption/production up to this moment. These widgets come from http://oilprice.com, a very intersting side which has many intelligent comments. I do not know who is behind the site, and I am not paid anything for the link :-)|
|10 Sep 11||1.
Maintenance and recalibration of CO2 and O3 sensors|
2. Overall power consumption of the gas sensors and the air zero generator is less than 700 W, so a 1000VA UPS should suffice. I ordered a refurbished APC at www.battery-direct.de, an online shop I highly recommend for UPS replacement batteries (and second hand refurbished UPS, for the not so wealthy!). The aim is to replace the big blown (<100kg) 5000VA UPS by two or 3 smaller devices, which I will be able to carry without breaking my back.
|23 Aug 11||The malfunction porblem reappeared after a couple of hours. As I am abroad, I can not find the problem by remote diagnosis. I have restarted everything, and it seems ok. Be patient! In the worst case, there will not be live data for the next 10 days.|
|21 Aug 11||There was a problem with the automatic downloading form the logger which started the 19th Aug.. Things seem to be working again.|
|02 Aug 11||
Quick calibration of CO2 and O3 sensors. I am unable to bring
nback to life the NOx sensor AC31M going. So we will have to
live without these data for the next weeks.
I added a graph showing the asymptotic CO2 trends (and those of Mauna Loa) to the "trends" page.
These asymptotic CO2 values which would exist if wind velocity was infinite (and the atmospheric boundary layer well mixed up) are reasonably close to the Mauna Loa measurements, adjusted for the positive latitudinal gradient.
|01 Aug 11||
Very serious hardware failure after 31Jul11 18:30 UTC.
The big APC UPS died, blowing all main fuses.
The radioactivity computer does not boot anymore; the AC31 NOx sensor remains dead; a fault with the automatic powering ON of the logger computer was found.
I try my best to bring station up again, but will probably keep the NOx and radioactivity measurements shut off. More on this asap.
18:35: This was really a black day, and I am not sure that all problems have been solved.
On top of the disfunctions mentioned above, was a logger computer who did not restart properly after a power loss (cause: flat Bios battery) and a RS232 cable which had a broken line.
I managed to bring back to life everything except the NOx sensor, which might be Off for the rest of the months. There seems to be no large data loss. For the moment, the automatic transfers from the logger to the Linux server making the graphs seem ok again...
Upload jul11.dat to data archive. There are no missing data in this month.
|29 Jul 11||Maintenance by Envicontrol|
|01 Jul 11||Uploaded jun11.dat into data archive. Cleared logger memory.|
|25 Jun 11||
Maintenance and recalibration work. There were two serious
1. the backup Davis Vantage Pro console was blank, someone (cleaning staff?) had pulled out the power cord. The station didn't restart properly, even after making a cold boot. After much fiddling around, ir came back to life.
2. The pump which pumps distilled water into the WBGT device was blocked. Dismounting, checking...ok again.. Possibly the last 7-14 days wet bulb (and WBGT) data are incorrect. As the weather was awful, the consequences are not serious.
|17 Jun 11||The UVI from meteoLCD is used by the Belgian Institute of Space Aeronomy on their page with live UVI data from Belgium and Luxembourg: see here! I added the link at the live data page.|
|03 Jun 11||New paper by F. Massen on obtaining sunshine duration from pyranometer readings.|
|01 Jun 11||Uploaded may11.dat to data archive|
|24 May 11||Maintenance and calibration work.|
1. Previous day problem seems to be solved; may be an
overheating by the computer who collects data from the logger.
2. Prof. Roger Pielke Sr. asked me to write a guest post on DTR on his weblog. You may read it here.
|13 May 11||
There was a serious computer problem starting two days ago, no
new data were uploaded to the webserver.
Problem possibly solved now (21:32), but graphs may take some time to recover.
|01 May 11||Uploaded apr11.dat to data archive|
|29 Apr 11||Maintenance of gas sensors by Envicontrol|
|26 Apr 11||
1. The Peltier cooling/warming of the UVB sensor (ch.9) seems to
have gone down the 19Apr11 at 17:30. Since that time the thermal
regulation was practically inexistent. All UVB readings from
19Apr 17:30 to 26Apr 07:30 should be considered suspect (even if
the values seem ok at a first glance). The problem probably
comes from a bad contact in an outside connector.
2. WBGT sensor restarted at 08:00
|06 Apr 11||
1. Pyranometer/Ampbox problem seems definitively solved
2. Uploaded mar11.dat file to the data archive
|02 Apr 11||
1. Maintenance and recalibration work on gas sensors
2. The pyranometer problems scould to be solved (I changed a common GND connection); nightime readings seems stable now.
|21 Mar 11||There is a serious problem with the pyranometer system: nighttime readings vary from 8 to over 30 Wm-2! The problem surely comes from the new Kip & Zonen amplifier box. Some tests made this morning gave a stable black offset of about 8 Wm-2, the same as that displayed during last night. I mounted a second old pyranometer on terrasse, together with a new mV amplifier from Vernier. The data are read by a vintage Mikromec logger. This test assembly will run for 7 days.|
|07 Mar 11||The Kip&Zonen pyranometer (model CM21, ch.6 in the data file) is again on-line: I installed the new AMPBOX (an amplifier box with a 4-20mA output) and added a separate old pyranometer attached to a logger to make some checks over the next days.The datalogger was emptied and restarted with a slightly modified calibration function for the pyranometer ch.6 at 16:00|
|05 Mar 11||Uploaded feb11.dat file into data archive.|
|24 Feb 11||The NOx sensor again went bad during the 19th February. Probably this is again an ozone generator problem whose origin is unknown. I solved the previous breakdown by increasing the ventilation of the sensor (opening the whole enclosure to increase heat removal). Remedial action will not be possible before the 28th Feb.|
|19 Feb 11||Maintenance and calibration work.|
|08 Feb 11||Uploaded jan11.dat file to data archive.|
|02 Feb 11||
There is a serious problem with the NOx sensor: after working ok
for some time, its reading dive to zero.
I have not yet found the cause of the problem.
|Made overlay of 2010 total ozone column measurements at Uccle (KMI, Brussels, Belgium) and Diekirch). Also added animated gif with all meteoLCD WOUDC graphs from 2001 to 2010 and another two with the UCCLE Brewer #16 graphs from 1983 to 2010 and both meteoLCD and Uccle for 2001-2010.|
|22 Jan 11||
Rechecked calibration of CO2, O3 and NOx sensors; cleaned air sample inlet filter.
There were a couple of problems following the last Envicontrol maintenance:
1. Sample inlets of CO2 and O3 sensors were not connected. This does not have an important influence for CO2, but may lower O3 readings
2. NOx sensor was left in ppb mode, and not put back into ug/m3 mode. This means that NO readings from 14/01 to 22/01 08:15 are to low by a factor of 1.35, and NO2 too low by a factor of 2.05
|14 Jan 11||Maintenance and calibration work by Envicontrol. Discard the spikes!|
|09 Jan 11||
1. I am busy updating and recalculating the local trends
from 1998 to 2010. See the solar energy and air temperature
trends, which are in a beautiful agreement with the "fast
process sensitivity" found by Scafetta. As usual the very poorly
funded meteoLCD station with its volontary crew is the first
showing these trends for Luxembourg.
2. There remains a problem with the rain sensor graphics. Will probably be self-healing.
|08 Jan 11||
1. Rain sensor was blocked by bird droppings during
06 an 07 Jan 2011.Taking into account the accumulated water
volume in the receiving bucket and the cumulative rainfall
during the same period registered by the sensor (as blockage was
not 100%), total precipitation from 06Jan11 00:00 to
07Jan11 midnight would be 30.1mm. This is the period of
heavy rainfall leading to high water levels and some flooding of the river Sure
here in Diekirch. I replaced the collector of the sensor by
another one, which has a slightly bigger outlet and a more
2. CM21 Pyranometer problem found: the amplifier 20mA box (CT24, IP66) mounted in the open developed a leak and was full of water. Circuit board is corroded, a transistor lost one leg, no repair possible. I will try to order asap a replacement (the Ampbox from Kip&Zonen, replacing the disused CT24). Meanwhile solar irradiance data will continue to be fetched from the Vantage-Pro-Plus station.
|01 Jan 10||
1.The usual new year problem with the 7 days
graphs.. as the logger file does not contain the year, Gnuplot
gets confused the 1rst week of every new year. This will self
correct the 7th January.
2. Solar radiation (ch.6) still from Vantage Pro Plus backup station.